### v2

# GeoGuess: Multimodal Reasoning based on Hierarchy of Visual Information in Street View

Abstract
Multimodal reasoning is a process of understanding, integrating and inferring information across different data modalities. It has recently attracted surging academic attention as a benchmark for Artificial Intelligence (AI). Although there are various tasks for evaluating multimodal reasoning ability, they still have limitations. Reasoning on hierarchical visual clues at different levels of granularity, i.e., local details and global context, is of little discussion, despite its frequent involvement in human reasoning. To bridge the gap, we introduce a novel and challenging task for multimodal reasoning, namely GeoGuess. Given a street view image, the task is to identify its location and provide a detailed explanation. In this work, we establish a benchmark for GeoGuess by introducing a specially curated dataset GeoExplain which consists of panoramas-geocoordinates-explanation tuples. Additionally, we present a multimodal and multilevel reasoning method, namely SightSense which can make prediction and generate comprehensive explanation. Our analysis and experiments demonstrate their outstanding performance in GeoGuess.
1 Introduction
The eye sees only what the mind is prepared to comprehend. – Robertson Davies
Figure 1: An GeoExplain example: panoramas-geoordinates-explanation tuple. Each sample consists of a set of panoramas, geoordinates and explanation on how to infer the location. Hierarchical visual information, i.e., local details and global context are both necessary in this example.
Humans can not only see the world, but more importantly comprehend the world through it. Observing, associating visual or textual knowledge, making interpretation, and even building a cognitive framework demonstrate a critical and widely used ability of human intelligence, known as multimodal reasoning. Therefore, developing more human-like multimodal reasoning ability is an essential step towards advancing Artificial General Intelligence (AGI) [15].
In recent years, there has been significant enthusiasm for researching on Large Language Model (LLM) [49, 6, 11, 63] driven by the rise of self-attention mechanism and Transformer architecture [64]. We have witnessed the remarkable performance of LLM on Natrual Language Understanding [43, 51, 76]. The great progress prompted the AI community to embark on an exploration of Multimodal Large Language Model (MLLM) [50, 3, 30] and multimodal reasoning [4, 72] that incorporate visual context as additional input. Despite the success achieved by techniques such as the Chain-of-Thought (CoT) [71], Tree-of-Thoughts (ToT) [77], and Self-Consistency [70], some studies highlighted that state-of-the-art MLLMs still struggle in multimodal reasoning [13, 54].
In response, a diverse range of multimodal reasoning testbeds have been developed, such as Science Question Answering (ScienceQA) [41], Geometric Question Answering [7] and GQA [24]. Although these testbeds vary in domains and disciplines, we believe they still face the following limitations:
  • Superficial Reasoning Process: These testbeds were relatively simple, requiring only superficial reasoning process and limited external knowledge which cannot embody the diverse, hierarchical and complex nature of human reasoning. This limitation weakens the practical grounding of expert-level multimodal reasoning.
  • Insufficient Visual Context at Different Granularity: Existing multimodal reasoning testbeds primarily focus on reasoning over global visual context, often neglecting local visual clues. Identifying tiny visual clues and extracting meaningful information from them is also an intuitive aspect of human cognition. Lack of multi-level visual understanding restricts reasoning performance on detail-oriented tasks.
To this end, we introduce a new task for multimodal reasoning: GeoGuess. Given a street view image, the task is to predict the location depicted in the image and provide a comprehensive explanation. GeoGuess is a variant of image geo-localization task [22, 81]. Existing exploration of image geo-localization only requires predicting coordinates of the location in a black box manner. On the contrary, GeoGuess places greater emphasis on reasonable interpretations in natural language as illustrated in Figure 1. GeoGuess is one of the most complex and challenging tasks designed for multimodal reasoning. Participants must analyze hierarchical visual elements, including global context like landscape and local details like car plate. The complexity reflects the multifaceted nature of real-world reasoning, making it a challenging endeavor even for humans. Therefore, a system that excels in GeoGuess must identify tiny clues in the image, grasp the overall vibe, associate geographic knowledge, reason between these elements and draw a final inference.
Furthermore, we introduce a novel dataset, GeoExplain, collected from the image geo-localization game GeoGuessr [14]. Each sample in GeoExplain includes a set of panoramic images, the geo-coordinates of the location and an explanation of how the exact location can be inferred from the image contents. We remove locations that are easy to recognize to make GeoGuess an expert-level task even for humans. We also propose a three-stage method, SightSense, specially designed for hierarchical multimodal reasoning to leverage local and global visual understanding and external knowledge. More specifically, in the first stage, the model identifies fine-grained visual clues via an open-vocabulary object detection. The detected objects are then used to retrieve knowledge snippets in the second stage. Finally, these elements and global context are fused through an adapter, enabling the model to generate a comprehensive response. Our experiments demonstrates the effectiveness of SightSense on GeoGuess, outperforming other state-of-the-art baselines.
We summary this work’s contribution as follows:
  • We propose a new task for multimodal reasoning: GeoGuess. It involves a more complex reasoning process and more comprehensive understanding of the visual contents which brings new challenges in the development of MLLM and AGI.
  • We propose a novel dataset specifically designed for GeoGuess, GeoExplain. To our best knowledge, GeoExplain is the first dataset to provide natural language explanation of how a location can be referred from the image contents.
  • We propose a new three-stage method SightSense for hierarchical multimodal reasoning to identify visual hints from different levels of granularity, retrieve and leverage external knowledge.
  • We also conduct experiments on GeoExplain. The result demonstrated that our approach surpasses other baselines models.
2 Related Works
2.1 Image Geo-Localization
Recent years have witnessed remarkable advancements in the field of image geo-localization, a subdomain of computer vision. The task of image geo-localization involves estimating the precise geographic coordinates where the photograph was taken. Recent developments in image geo-localization can be categorized into three main approaches: retrieval-based method, regression-based method and LLM-based method.
The retrieval-based method [53, 57, 58, 62, 82] searches for the most similar photo within a large-scale photo set in which each photo is annotated with its exact geographical coordinates. The coordinates of the best-matched image are then identified as the result. The regression-based method [19, 66, 47, 55, 67, 73] approaches the geo-localization task as an earth segments regression problem, employing deep learning networks in a black-box manner. Recent LLM-based method [38, 21, 75, 31] utilize the reasoning ability obtained from large-scale training corpus to generate prediction in natural language.
Despite the significant success achieved by these methods, each approach still faces notable challenges. For retrieval-based method, matching images captured from different viewpoints at the same location remains difficult, and the approach demands substantial computational resource due to searching in an additional large database. Besides, some local clues may be overlooked because they may be disappeared in different photos even at the same location. Regression-based method suffers from lack of interpretability. As the regression process functions in a black-box manner, it is difficult to understand how predictions are made, making the results less transparent and less persuasive. Neither approach follows the human way of thinking, which involves sensing the global vibe from a photo, such as terrain or landscape, identifying subtle details, such as plants or animals, and connecting them with geographical knowledge to infer the exact location. On the other hand, existing LLM-based method suffers from poor performance on non-landmark scene and hallucination [23].
2.2 Visual Language Modelling
Scaling up the size of models’ parameters and training data has significantly enhanced the capabilities of attention-based Large Language Models (LLMs) [49, 6, 11], enabling them to surpass human performance in various generative tasks in Natural Language Processing (NLP), such as question answering [48, 8] and text summarization [39, 79]. However, most of them were limited by incapacity of reading visual information. Conversely, Large Vision Model (LVMs) [12, 40, 18] can see clearly but cannot talk. Building on the success of both LLM and LVM, significant research efforts have been dedicated to extending the impressive zero-shot capabilities of LLMs to the visual domain by incorporating visual modalities of LVM into generative tasks, which is called Multimodal Large Language Model (MLLM), such as GPT-4 [1], LLaVA [35] and Llama [17]. An MLLM can bridge effective connections between different gaps, thus succeeding in various multimodal tasks, including 1) image-grounded question-answering dialogue system [10, 44, 45], 2) image-grounded chit-chat dialogue system [59, 60], 3) visual-evidence-embedded dialogue system [56, 32], 4) image-response dialogue system [78, 61], 5) video-grounded question-answering dialogue system [26, 2, 27], 6) video-grounded chit-chat dialogue system [20, 34, 9] and 7) video-grounded real-time comment dialogue [42, 69].
3 GeoExplain
To craft a dataset specifically for GeoGuess Task is evitable due to lack of similar training corpus. To this end, we collected a brand new dataset, GeoExplain, which consists of 5459 panorama-geoordination-explanation tuples, represented by Equation 1.
$G=\{(\mathbf{P_{i}},L_{i},E_{i})\}_{i=1}^{n}$ (1)
Each tuple corresponds to a specific geographic location and $n$ is the total number of samples. The panorama $\mathbf{P_{i}}=\{I_{1},I_{2},\dots,I_{im}\}$ is a series of street view images of the location from different angles, $L_{i}$ is the exact location on street level, and the explanation $E_{i}$ provides the expert-level natural-language description how human can infer the location from the visual information in $\mathbf{P_{i}}$. Figure 1 demonstrates an example of GeoExplain. We collected GeoExplain from an image geo-localization game GeoGuessr and its community on Reddit [52].
In this section, we firstly introduce the data collection method and then present an analysis of our dataset. To our best knowledge, GeoExplain is the first dataset which includes complete textual explanations how a location can be inferred from a panorama on expert level.
3.1 GeoKnowledge
To incorporate external geographic knowledge, we collect a domain-specific knowledge base, GeoKnowledge. GeoKnowledge consists of image-snippets pairs indexed by country where the snippets tell how to identify a country from visual clues in the image. GeoKnowledge can be represented by Equation 2, where $KI_{j}$ is an image of country-specific object and $S_{j}$ is the corresponding knowledge snippets.
$K=\{(KI_{j},S_{j})\}_{j=1}^{k}$ (2)
3.2 Data Collection
GeoExplain is collected from an online geographical game GeoGuessr and its community on Reddit. GeoGuessr challenges players to guess locations around the world based on Google Street View images. The game offers daily challenge mode where players are given 5 locations to guess. Thanks to its large player base and active community on Reddit, players who accurately infer these locations share their expert-level reasoning process in the daily challenge discussion threads on Reddit. We firstly selected 6085 locations from 1217 historical daily challenge from June 2021 to Sep 2024. We manually filtered these locations by following steps.
  • 1) We firstly remove locations that are close to landmarks or have clearly visible landmarks, e.g., Eiffel Tower in Paris or Opera House in Sydney.
  • 2) We then filtered out locations where no reasonable and well-supported explanations of how to find the place were posted in the corresponding discussion threads.
  • 3) We also filtered out high-frequency places and ensure they are distributed across different continents, landscapes and both urban and rural areas.
After carefully filtering locations, we collected street view images in different angles from Google Street Views [16]. On the other hand, we extracted corresponding expert-level reasoning explanations from Reddit discussions. Specifically, posters usually include their scores in the discussion, so we first splited the game scores and reasoning processes. Then, we filtered these explanations as follows: 1) Remove all explanations with scores lower than 4000 (with max 5000), 2) Remove all explanations that are not in English, 3) Remove all explanations that rely on non-reasoning factors to arrive at the answer, such as living nearby or having previously finished the same challenge.
GeoKnowledge is collected from a geographical knowledge website [^1] Plonk It which is established specially for sloving the image-geolocalization puzzles.
3.3 Data Analysis
We present an analysis of GeoExplain in this section. There are in total 5459 locations, 169972 panorama images and 41068 explanations on how to find the location based on images. We endeavor to ensure the richness and diversity of vocabulary in these explanations. The locations in GeoExplain are distributed across all continents, covering a wide range of terrains. They include diverse landscapes such as mountains, deserts, forests, coastlines, and urban environments, ensuring a comprehensive geographical representation. Regarding panorama statistics, each location contains an average of 31.13 street view image snippets, with a maximum 33 and a minimum 3. Their size vary, ranging from (600 * 280) to (512 * 512). On the other hand, GeoKnowledge covers 134 countries or regions, encompassing all those included in GeoExplain.
To measure GeoExplain’s difficulty of questions, reliance on reasoning and request for visual clues at different levels of granularity, we conducted a survey between our dataset GeoExplain and other two similar multimodal reasoning datasets: ScienceQA and GeoQA. Specifically, we randomly sampled 100 examples from GeoExplain, ScienceQA and GeoQA respectively. Volunteers were asked to rate them on a scale from 0 to 2 in following questions. In order to avoid the bias between raters, we anonymized all samples and averaged ratings from three volunteers. Since, even after anonymization, peers could still easily identify which dataset a sample came from, we randomized the order of all samples and invited three non-peers with no prior knowledge of the basic information of these three datasets to rate them.
  • The level whether the answer needs complex reasoning process;
  • The level whether the answer requires both local visual clues and global context;
  • The level whether the answer is reasonable and understandable without external knowledge;
  • The level of difficulty that the answer can be correctly deduced.
Figure 2: Human Analysis on GeoExplain and Other Multimodal Reasoning Dataset
From Figure 2, GeoExplain demonstrate the highest level in all four aspects. The results indicate GeoGuess is a task that ivolves more complex reasoning, at expert-level difficulty, even for humans.
4 SightSense
Figure 3: Architecture of SightSense: The first stage is Visual Clue Detection which detects identifiable visual clues. The second stage is Multimodal Knowledge Retrieval to incorporate external geographical knowledge. The final stage is to concatenate visual clues, global context and knowledge snippets as prompt to generate reasoning.
To tackle GeoGuess, We propose SightSense, a novel multimodal reasoning approach that integrates hierarchical visual information across multiple levels of granularity including local clues and global context. More specifically, SightSense captures fine-grained visual clues, perceives the coarse-grained landscape, and leverages external knowledge to support reasoning. In this section, we elaborate on the design details of SightSense.
4.1 Problem Statement
We formulate SightSense task as follows. Suppose we have a GeoExplain set $G=\{(\mathbf{P_{i}},L_{i},E_{i})\}_{i=1}^{n}$ and another GeoKnowledge base $K=\{(KI_{j},S_{j})\}_{j=1}^{k}$ as what we mentioned in Section 3. The goal of SightSense is to learn a generative model $P(L,E|P,K;\theta)$ from $D$ where $\theta$ denotes the parameters of the model. Therefore, we can generate location and corresponding explanation according to Equation 3.
$L^{*},E^{*}=\arg\max_{L,E}P(L,E|P,K;\theta)$ (3)
4.2 Model Architecture
In order to fully leverage visual information at different levels of granularity, SightSense uses a three-stage method. Figure 3 illustrates the overall workflow. The three stages are Visual Clues Detection, Multimodal Knowledge Retrieval and Reasoning Generation respectively. In the first stage, potentially identifiable local elements are extracted from street view images $\mathbf{P}$ by using an open-vocabulary object detector. The second stage retrieves relevant geographical knowledge $KI,S$ based on these visual clues. Finally, the reasoning stage integrates global context, detected visual clues and retrieved external knowledge to generate informative answer and explanation L,E{L,E}.
4.3 Visual Clue Detection
To capture and integrate fine-grained visual details into the reasoning process, in the first stage, we explicitly extract all possible identifiable fine-grained elements $\mathbf{V_{i}}=\{V_{1},V_{2},\dots,V_{in}\}$ by an open-vocabulary object detector according to the Equation 4. Each element $V_{ij}$ is a cropped image fragment from original panorama $\mathbf{P_{i}}$, containing one or more specific potentially identifiable local elements.
$\mathbf{V_{i}}$ $=\text{OOD}(\mathbf{P_{i}},prompt)$
$=\cup_{j=1}^{im}\text{OOD}(I_{j},prompt)$ (4)
Based on analysis of high-frequency nouns in the explanation $E$ and knowledge snippets $S$, we design a comprehensive prompt to capture all potentially identifiable object. We use Grounding DINO[37] as our open-vocabulary object detector in this stage.
4.4 Multimodal Knowledge Retrieval
Since GeoGuess is an expert-level task even for humans, relying on the inherent commonsense knowledge in MLLMs obtained from large-scale pretraining, or on reasoning techniques like [71, 77], is insufficient for generating accurate answers. To alleviate potential hallucinations produced by LLM, we have to explicitly incorporate external geographical knowledge into the reasoning process. Accordingly, in the second stage, we aim to retrieve informative knowledge snippets from GeoKnowledge in a multimodal way to assist the reasoning process, inspired by [28].
More specifically, based on the visual clues $\mathbf{V_{i}}$ detected in the first stage, we retrieve top 3 visually similar country-specific object images in GeoKnowledge and exploit their knowledge snippets as assisting external knowledge. Meanwhile, corresponding potentially identifiable objects are chosen as assisting local clues to participate in the reasoning process. The retrieval process can be formulated as Equation 7 where $\mathbf{S^{*}}$ are assisting knowledge snippets and $\mathbf{V^{*}}$ are assisting local clues.
$\mathbf{S^{*}}$ $=\{S_{i_{1}^{*}},S_{i_{2}^{*}},S_{i_{3}^{*}}\}$ (5)
$\mathbf{V^{*}}$ $=\{V_{j_{1}^{*}},V_{j_{2}^{*}},V_{j_{3}^{*}}\}$ (6)
$i_{1}^{*},i_{2}^{*},i_{3}^{*},j_{1}^{*},j_{2}^{*},j_{3}^{*}$ $=\arg\operatorname{Top\text{-}3}_{i,j}\frac{\text{VE}(KI_{i})\cdot\text{VE}(V_{j})}{|\text{VE}(KI_{i})||\text{VE}(V_{j})|}$ (7)
In Equation 7, $KI$ denotes an image of country-specific object from GeoKnowledge, and $V$ represents detected visual clues obtained in the first stage. To retrieve the top 3 similar image in GeoKnowledge, we employ cosine similarity. A visual encoder, denoted as VE, is used to embed images into vector representations. In our work, we adopt CLIP-ViT [50] as the visual encoder with an embedding dimension of 768.
4.5 Reasoning Generation
In the reasoning generation stage, we integrate hierarchical visual information and external geographical knowledge. There are two types of input in a sequence. More specifically, for a specific sample, we first concatenate the global context, i.e., the thumbnail image of the panoramas, and the image of assisting local clues to construct the image prompt. Then we use retrieved knowledge snippets to construct the text prompt.
In text prompt, we explicitly instruct the MLLM to reason based on provided global context, local clues and relevant external knowledge. We design a standard answer format: ”PLACE {COUNTRY, STATE, CITY, STREET}. EXPLANATION.” This ensures that models can generate well-structured location prediction for easier evaluation. Even when using un-finetuned baseline models in experiments, the prompt can instruct the model to attempt to generate an explanation.
After constructing both image and text prompt, we feed them into a pre-trained MLLM to perform reasoning and generate final answer.
4.6 Training Policy
We adopt a straightforward autoregressive language modeling objective by optimizing Cross-Entropy Loss. The training objective can be formulated as Equation 8 where $x_{t}$ denotes the ground truth token at position $t$.
$\text{Loss}=-\sum_{t=1}^{T}P(x_{t}|x_{1},\dots,x_{t-1})$ (8)
5 Experiments
We conduct comprehensive experiments comparing SightSense with state-of-the-art baselines on GeoExplain dataset to evaluate their performance on the task GeoGuess. These experiments allows us to investigate the difficulty of GeoGuess and the advantages of SightSense in integrating hierarchical visual information for improving reasoning. In this section, we elaborate on our experimental setup and analyze the results through both quantitative metrics and qualitative case studies.
5.1 Experiments Setup
SightSense utilizes GroundingDINO [37] as the open-vocabulary object detector in the visual clue detection stage, CLIP-ViT [50] as the visual encoder in the multimodal knowledge retrieval stage and Qwen2VL-7B-Instruct [68] as the base reasoning model in reasoning generation stage. We finetune only parameters in Qwen2VL and other baselines, while keeping the parameters of models in GroundingDINO and CLIP-ViT frozen. For finetuning details, we adopt Adam optimizer [25] to optimize the training objective and use LinearLR to adjust the learning rate. All training work was completed on NVIDIA H100 PCIe.
5.2 Evaluation Metrics
We refer to the location and explanations generated by SightSense and other baselines as candidate location and explanations. Evaluation of the GeoGuess task comprises two aspects: (1) the accuracy of the candidate location, and (2) the textual similarity between model-generated candidate explanations and human-written references which is measured by both automatic metrics and human evaluation.
5.2.1 Accuracy of Location
We convert the ground-truth geo-coordinate into structured ground-truth location in the format (COUNTRY, STATE, CITY, STREET) which matches the format of candidate location. In case that some countries may have different names, we apply fuzzy match [^2] RapidFuzz and pycountry. Accuracy is computed at each of the four hierarchical levels as Equation 9, where Match is the fuzzy match function, $c_{i}$ and $gt_{i}$ denotes the candidate location and ground-truth location respectively.
$\text{Acc}=\frac{\sum_{i=1}^{n}\text{Match}(c_{i},gt_{i})}{n}$ (9)
5.2.2 Explanation: Automatic Metrics
To measure the similarity between candidate and reference explanations, we employ two aspects of automatic metrics.
  • Relevance: We use (1) BLEU [46], (2) Rogue-L [33], (3) CIDEr [65], (4) Meteor [5] and (5) BERTScore [80] to evaluate the relevance between candidate and reference explanations.
  • Diversity: We use (1) Dist-1 and (2) Dist-2 [29] to measure the diversity of candidate explanations themselves.
Type Methods BLEU_3 BLEU_4 ROUGE_L CIDEr METEOR BERTScore distinct_1 distinct_2
P
LLaVA-1.5-7B 2.99 0.91 18.55 1.17 28.84 84.85 6.08 21.34
GPT-4o-mini 1.53 0.00 13.95 0.22 19.44 82.59 5.32 26.30
DeepSeek-VL2-tiny 0.07 0.00 1.34 0.06 6.68 79.63 14.51 32.54
smileGeo 0.58 0.00 7.57 0.08 10.15 79.97 12.24 28.55
GeoReasoner 1.39 0.09 9.19 0.15 17.87 81.33 7.86 25.80
SightSense_wKVC (ours) 3.71 3.92 21.13 5.51 22.71 84.46 8.57 18.45
P+VC
LLaVA-1.5-7B-VC 3.00 $\uparrow$ 0.90 $\downarrow$ 18.85 $\uparrow$ 1.03 $\downarrow$ 28.29 $\downarrow$ 84.98 $\uparrow$ 5.02 $\downarrow$ 17.70 $\downarrow$
GPT-4o-mini-VC 2.12 $\uparrow$ 0.56 $\uparrow$ 12.62 $\downarrow$ 0.33 $\uparrow$ 23.76 $\uparrow$ 82.88 $\uparrow$ 6.92 $\uparrow$ 28.90 $\uparrow$
DeepSeek-VL2-tiny-VC 0.11 $\uparrow$ 0.04 $\uparrow$ 1.75 $\uparrow$ 0.26 $\uparrow$ 9.11 $\uparrow$ 79.79 $\uparrow$ 18.29 $\uparrow$ 35.32 $\uparrow$
SightSense_wK (ours) 6.18 $\uparrow$ 4.18 $\uparrow$ 26.88 $\uparrow$ 6.82 $\uparrow$ 29.81 $\uparrow$ 86.18 $\uparrow$ 9.39 $\uparrow$ 19.50 $\uparrow$
P+VC+K
LLaVA-1.5-7B-KVC 5.44 $\uparrow$ 2.03 $\uparrow$ 20.00 $\uparrow$ 1.12 $\uparrow$ 29.35 $\uparrow$ 85.39 $\uparrow$ 4.58 $\downarrow$ 15.37 $\downarrow$
GPT-4o-mini-KVC 2.45 $\uparrow$ 0.40 $\downarrow$ 13.32 $\uparrow$ 0.37 $\uparrow$ 22.42 $\downarrow$ 82.88 - 7.49 $\uparrow$ 28.16 $\downarrow$
DeepSeek-VL2-tiny-KVC 1.16 $\uparrow$ 0.27 $\uparrow$ 5.46 $\uparrow$ 0.56 $\uparrow$ 12.47 $\uparrow$ 83.36 $\uparrow$ 13.49 $\downarrow$ 33.87 $\downarrow$
SightSense (ours) 8.93 $\uparrow$ 5.09 $\uparrow$ 31.58 $\uparrow$ 10.43 $\uparrow$ 36.66 $\uparrow$ 87.71 $\uparrow$ 9.00 $\downarrow$ 19.23 $\downarrow$
Table 1: Automatic evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines and underlying number denotes the highest accracy for each type. $\uparrow$, $\downarrow$ and - indicates how the accuracy change for corresponding models when introducing more modules.
5.2.3 Explanation: Human Evaluation
Considering the complexity involved in GeoGuess explanation generation, automatic metrics may be insufficient for reliably evaluating reasoning quality and the adequacy of supporting evidence. Therefore, human evaluation is adopted for a more comprehensive analysis. More specifically, we randomly select 100 samples from the test set of GeoExplain. Three voluntary experts in GeoGuessr are asked to rate candidates generated by SightSense and other baselines on a scale from 0 to 2 in following questions.
  • Fluency: whether the explanation is fluent and natural.
  • Evidence: whether identifiable visual evidence is explicitly mentioned in the explanation.
  • Knowledge: whether the explanation incorporates relevant external geographical knowledge or commonsense.
  • Reasoning: whether the reasoning process connect the evidence with external knowledge to generate a logically sound, correct, and understandable explanation.
  • Overall Quality: overall quality of the generated explanation.
5.3 Baselines
We compare SightSense with several state-of-the-art baselines in multimodal reasoning. We validate the effectiveness of hierarchical visual information in the reasoning process even for other baselines by transplanting the first two stages of our pipeline into them. This enables a fair comparison and highlights the generalizability of our approach. Furthermore, we classify the baselines into three different types according to information they receive during generating explanations. Note we do not finetune on models with * due to their close-source nature or inapplicable training policy.
•  Panorama Only: In this type, we use SightSense_wKVC which is a simplified version of SightSense, removing the Visual Clue Detection and Knowledge Retrieval. We selected both geo-localization specified methods and well-established and widely recognized state-of-the-art MLLMs with comparable model sizes: (1) GeoCLIP [66]; (2) StreetCLIP [19]; (3) smileGeo [21]; (4) GeoReasoner [31]; (5) LLaVA-1.5-7B [36]; (6) GPT-4o-mini* and (7) DeepSeek-VL2-tiny [74]. These models serve as strong and representative baselines for evaluating the performance of our approach.
•  Panorama + Visual Clue: Similarly, we finetune SightSense_wK which removes only the Knowledge Retrieval from SightSense. Meanwhile, we transplant the Visual CLue Detection stage into other MLLM baselines. We denote them as (1) LLaVA-1.5-7B-VC; (2)GPT-4o-mini-VC* and (3) DeepSeek-VL2-tiny-VC.
•  Panorama + Visual Clue + Knowledge: We use SightSense with all components enabled. Accordingly, we refer to the baseline models with all three modules as (1) LLaVA-1.5-7B-KVC; (2) GPT-4o-mini-KVC* and (3) DeepSeek-VL2-tiny-KVC.
5.4 Main Results
We summarizes results of location accuracy in Table 2, automatic metrics for explanation in Table 1 and human evaluation for explanation in Table 3.
Type Methods Country State City Street
P
LLaVA-1.5-7B 16.80 0.37 0.00 0.00
GPT-4o-mini 4.58 0.55 0.00 0.00
DeepSeek-VL2-tiny 9.17 1.10 0.09 0.00
GeoCLIP 55.16 16.89 3.38 1.14
StreetCLIP 48.75 7.38 2.36 0.81
smileGeo 23.44 2.27 1.52 0.14
GeoReasoner 25.81 2.46 1.13 0.54
SightSense_wKVC (ours) 30.08 6.68 1.60 0.09
P+VC
LLaVA-1.5-7B-VC 16.62 $\downarrow$ 1.38 $\uparrow$ 0.00 - 0.00 -
GPT-4o-mini-VC 11.38 $\uparrow$ 0.92 $\uparrow$ 0.00 - 0.00 -
DeepSeek-VL2-tiny-VC 29.79 $\uparrow$ 5.32 $\uparrow$ 0.09 - 0.00 -
SightSense_wK (ours) 33.46 $\uparrow$ 5.27 $\downarrow$ 2.41 $\uparrow$ 0.51 $\uparrow$
P+VC+K
LLaVA-1.5-7B-KVC 25.80 $\uparrow$ 4.91 $\uparrow$ 0.10 $\uparrow$ 0.09 $\uparrow$
GPT-4o-mini-KVC 12.37 $\uparrow$ 1.01 $\uparrow$ 0.12 $\uparrow$ 0.09 $\uparrow$
DeepSeek-VL2-tiny-KVC 32.36 $\uparrow$ 4.58 $\downarrow$ 0.37 $\uparrow$ 0.12 $\uparrow$
SightSense (ours) 60.97 $\uparrow$ 23.88 $\uparrow$ 6.11 $\uparrow$ 4.55 $\uparrow$
Table 2: Accuracy of candidate location in %. Denotations are the same as in Table 1
Method Flency Evidence Knowledge Reasoning Overall
LLaVA-1.5-7B 1.01 1.05 0.77 0.61 0.67
GPT4o-mini 1.34 0.68 1.08 0.21 0.52
DeepSeek-VL2-tiny 0.89 0.03 0.01 0.03 0.1
smileGeo 1.05 1.13 1.07 0.89 1.10
GeoReasoner 0.97 0.88 1.01 0.91 1.03
SightSense (ours) 1.16 1.59 1.67 1.11 1.58
Table 3: Human Evaluation on candidate explanation. Bold number denotes the highest accuracy among all baselines.
Figure 4: A case study of explanation generated by SightSense, LLaVA-1.5-7B, GPT-4o-mini and DeepSeek-VL2-tiny. Given image is a thumbnail of panoramas set and the groundtruth is on the right hand side.
•  Accuracy of Candidate Location: From Table 2, SightSense substantially outperforms all other baselines across all four levels, achieving notable gains of 27% and 18% on the Country and State levels respectively. Meanwhile, within each modality type, the simplified version of SightSense consistently achieves near-best performance except in the Panorama-only type, where SightSense_wKVC underperforms geo-localization-specified models. However, when all components are enabled, SightSense outperforms even these non-explanation geo-localization-specified methods.This verifies the effectiveness of the architecture and training objective of SightSense and highlights its advantage in addressing the GeoGuess task. The slight accuracy drop is due to misleading of confusing visual clues and potentially wrong retrieved knowledge snippets. We observe some illogical inconsistencies happens, such as accuracy drops at state level while they increases at city level. This suggests that finetuning may cause the model to confuse different administrative tiers. In general, the consistent accuracy gains across all four models further verify the robustness and generalizability of our method.
•  Automatic Metric of Candidate Explanation: We observe SightSense series still achieves the best performance regarding explanation similarity evaluation across all types in 1. Explanations generated by SightSense show the highest similarity with the reference explanations. The diversity metric is primarily dominated by DeepSeek-VL2 which means SightSense tends to generate a fixed template learned from the GeoExplain dataset. As with location accuracy analysis, we also notice significant increase in similarity metric when introducing Visual Clue Detection and Multimodal Knowledge Retrieval module, proving the effectiveness of our approach.
•  Human Evaluation of Candidate Explanation: Table 3 demonstrates the result of human evaluation. Except for Fluency, SightSense achieves the highest score across all other metrics. This indicates that the explanation generated by SightSense provide the strongest evidence, the broadest and the most relevant geographical knowledge, and the most complete reasoning chain. These improvements are achieved by the understanding and integration of hierarchical visual information and external knowledge proposed in SightSense.
5.5 Ablation Study
To better understand the contribution of each stage in SightSense, we provide a brief ablation study. From the comparison across the same model in different types, introducing Visual Clue Detection module results in 8 accuracy increases, 2 decreases and 6 ties. Similarly, introducing Multimodal Knowledge Retrieval module leads to 15 increases and 1 decrease. The comparison demonstrates that integrating hierarchical visual information and external knowledge consistently improves location prediction accuracy.
5.6 Case Study
To further investigate the performance of SightSense and other baseline models, we present a qualitative case study. As shown in Figure 4, SightSense is able to detect the fine-grained visual clues, i.e., Michigan license plate, and perceive the coarse-grained landscape, i.e., abundance of lakes. It then connect such hierarchical visual information with external knowledge to generate the final prediction which is accurate at both country and state level. In contrast, other baselines demonstrate notable limitation. GPT-4o-mini and DeepSeek generate significant hallucination that are inconsistent with visual evidence. LLaVA captures informative clues from the global context but fails to identify detailed visual elements, resulting in a vague location prediction.
6 Conclusion
In this paper, we introduce a novel multimodal reasoning task, GeoGuess, identifing the location of a street view and providing comprehensive explanation. To support GeoGuess, we construct a new dataset GeoExplain. To our best knowledge, GeoExplain is the first dataset that introduces expert-level explanation on how to infer the location from a given street view image. We further propose a three-stage method SightSense which understand hierarchical visual information at multiple levels to solve GeoGuess. Experiments on GeoGuess demonstrates our method outperforms other state-of-the-art baseline models and provide strong generalizability when transplanting to other models. Moreover, no existing model reach human-level accuracy, which validates the difficulty of GeoGuess and its significance as a benchmark toward the development of Artificial General Intelligence. Future work could focus on alleviating negative effects of misleading knowledge snippets to further achieve better performance on GeoGuess.